Solve the following system by substitution
8x-4y=16
y=2x-4
Show your solution below:
What is the solution for this system?
What type of system is this?
please help

Answer by stanbon(75887)

Solve the following system by substitution
8x-4y=16
y=2x-4
----------
Simplify the 1st equation:
2x-y = 4
Substitute to get:
2x-(2x-4)=4
2x-2x+4=4
4=4
=========
This means the two original equations are really the same;
one is a multiple of the other.
The system is described as "dependent".
The solution set is the set of all points on the line y = 2x-4
---------------------------
Cheers,
Stan H.